# Visuospatial navigation without distance, prediction, integration, or maps

### v3

@@ 4 Methods
How the elliptical manifolds, explicit for the indirect sequential agents (Fig. 2a) and implicit for the biased diffusive agents (Fig. S6), can be examined directly via perturbations. Perceptual, environmental, and movement related parameters all affect the shape and location of the manifolds (Fig. S4).
Inferring the generative mechanism, given the ellipses emanate from two adjacent corners, the manifolds may be marked by two raycasts simultaneously intersecting both corner locations, representing a change in wall perception in both corners at once. A numerical search for these ideal locations shows that simultaneous dual corner detection reconstructs elliptical arcs, with orientation rotated about the center between the two corners, and both perturbed FOV and visual resolution shifting the arcs predictably (Fig. S5 1st & 4th rows).
While perceptual and environmental parameters are fixed for a given training period, movement dynamics can shift manifold positioning as needed. Generated ideal elliptical arcs intersect with patch position at FOV of 40%, but not for 35 or 45. Though when trained with these FOVs, the agents learn to approximate ideal arcs useful for calculating patch position regardless (Fig. S5 2nd row). By increasing turning speed before intersecting with a manifold, the agent expands its capacity to perceive a corner, thereby allowing the agent to effectively precede ideal arc location. Looking at turning speed for each agent, it is clear that the agents with FOV 35 or 35% increase incoming turning speed in order to shift the ellipse to a more useful position (Fig. S5 3rd row). Thus, while various perceptual, action, and environmental parameters constrain elliptical manifold construction, the learning algorithm coordinates movement behavior such that generated ellipses facilitate navigational success.
Furthermore, mean field analysis can demonstrate how the two angle-based classes use the elliptical manifolds differently (Fig. S6). Indirect sequential agents (a) can sufficiently navigate without minima overlapping patch location since their algorithm relies on a sparse, stable subsample of orientations. Biased diffusive agents (b) navigate with respect to a disperse, dynamic sample of every orientation, thus their success depends on correlating mean field minima with patch location. Elliptical manifolds visibly structure attractors for both angle-based classes, and are not apparent with the smoother gradients of distance-based agents (c). Higher visual resolution angle-based individuals (bottom row, a/b) tend to exhibit increased elliptical manifold presence (south & east walls), possibly explaining why biased diffusive agents perform better at greater visual resolution.
@@
All movement correlations were transformations of the test trajectory data previously described. The first 25 timesteps, when the agent is orienting itself upon random initialization, is masked from movement statistics, the end of which represents an initial route heading.
The spatial heading profiles of the top left correlation plots in Fig. 2 were produced by plotting agent orientation relative to the initial route heading against the distance from the patch center. The polar histograms count the frequency of relative orientations, where the area under each bar is scaled to relative frequency.
The temporal correlation metric uses a function borrowed from a sperm chemotaxis paper [65], given the similarity between movement trajectory patterns:
@@
with $\phi$ representing the orientation angle of the agent and $t_{0}$ defined at a delay of 25 timesteps from initiation. Decorrelation time is defined by the length of time needed for $C{t}$ to drop below half.
Directedness, the spatial correlation metric, is based on the range of directional possibilities available to the agent, mapped across space. We calculate the information entropy [66, 67] of the orientational frequency distributions for each spatial bin across the environment:
@@
While the clusters are visually separable, unsupervised clustering algorithms learn to draw in a similar way (Fig. S9 bottom row). Although unsupervised clustering algorithms can designate boundaries, hybrid strategies between or far from cluster centers require qualitative context.

### v4

@@ 4 Materials and methods
How the elliptical manifolds, explicit for the indirect sequential agents (Fig 2a) and implicit for the biased diffusive agents (S6 Fig), can be examined directly via perturbations. Perceptual, environmental, and movement related parameters all affect the shape and location of the manifolds (S4 Fig).
Inferring the generative mechanism, given the ellipses emanate from two adjacent corners, the manifolds may be marked by two raycasts simultaneously intersecting both corner locations, representing a change in wall perception in both corners at once. A numerical search for these ideal locations shows that simultaneous dual corner detection reconstructs elliptical arcs, with orientation rotated about the center between the two corners, and both perturbed FOV and visual resolution shifting the arcs predictably (S5 Fig, 1st & 4th rows).
While perceptual and environmental parameters are fixed for a given training period, movement dynamics can shift manifold positioning as needed. Generated ideal elliptical arcs intersect with patch position at FOV of 40%, but not for 35 or 45. Though when trained with these FOVs, the agents learn to approximate ideal arcs useful for calculating patch position regardless (S5 Fig, 2nd row). By increasing turning speed before intersecting with a manifold, the agent expands its capacity to perceive a corner, thereby allowing the agent to effectively precede ideal arc location. Looking at turning speed for each agent, it is clear that the agents with FOV 35 or 35% increase incoming turning speed in order to shift the ellipse to a more useful position (S5 Fig, 3rd row). Thus, although various perceptual, action, and environmental parameters constrain elliptical manifold construction, the learning algorithm coordinates movement behavior such that generated ellipses facilitate navigational success.
Furthermore, mean field analysis can demonstrate how the two angle-based classes use the elliptical manifolds differently (S6 Fig). Indirect sequential agents (a) can sufficiently navigate without minima overlapping patch location since their algorithm relies on a sparse, stable subsample of orientations. Biased diffusive agents (b) navigate with respect to a disperse, dynamic sample of every orientation, thus their success depends on correlating mean field minima with patch location. Elliptical manifolds visibly structure attractors for both angle-based classes, and are not apparent with the smoother gradients of distance-based agents (c). Higher visual resolution angle-based individuals (bottom row, a/b) tend to exhibit increased elliptical manifold presence (south & east walls), possibly explaining why biased diffusive agents perform better at greater visual resolution.
@@
All movement correlations were transformations of the test trajectory data previously described. The first 25 timesteps, when the agent is orienting itself upon random initialization, are cut out when calculating movement statistics. Thus, the orientaion at t=25 represents the initial route heading.
The spatial heading profiles of the top left correlation plots in Fig 2 were produced by plotting agent orientation relative to the initial route heading against the distance from the patch center. The polar histograms count the frequency of relative orientations, where the area under each bar is scaled to relative frequency.
The temporal correlation metric uses a function borrowed from a sperm chemotaxis paper armonTestingHumanSperm2012, given the similarity between movement trajectory patterns:
@@
with $\phi$ representing the orientation angle of the agent and $t_{0}$ defined at a delay of 25 timesteps from initiation. Decorrelation time is defined by the length of time needed for CtC{t} to drop below half.
Directedness, the spatial correlation metric, is based on the range of directional possibilities available to the agent, mapped across space. We calculate the information entropy shannonMathematicalTheoryCommunication1949; pilkiewiczDecodingCollectiveCommunications2020 of the orientational frequency distributions for each spatial bin across the environment:
@@
Neural representation
The activation profiles of perceptron nodes were plotted with respect to spatial position in the environment and orientation, or head direction as head and body are not separable in our model. Datasets used for the grid plots (Fig 3A-C), same as for S6 Fig, consisted of node outputs for a static, uniform sample of agent locations and orientations. Datasets for the trajectory plots (Fig 3D-F) were the same as used for the movement correlations. Using either dataset, output activity was sorted via 50x50 spatial bins, averaged within each bin, smoothed with a Gaussian filter of standard deviation 1, and plotted from zero to maximum activation. The data was further sorted into 8 orientation bins and separately plotted. This procedure follows a recent paper that focused on constructed activation rate maps from neural networks trained to navigate vijayabaskaranNavigationTaskAction2022. Neural tuning curves were plotted by binning activation according to the angle between the agent and the goal, as in sarelVectorialRepresentationSpatial2017.
@@
While the clusters are visually separable, unsupervised clustering algorithms learn to draw in a similar way (S9 Fig, bottom row). Although unsupervised clustering algorithms can designate boundaries, hybrid strategies between or far from cluster centers require qualitative context.
Bottom-up dimensionality reduction
To complement the classification procedure, nonlinear dimensionality reduction algorithms provide a bottom-up, however less explainable means to separate navigational strategy. The set of actions related to all possible visual observations throughout the environment fully describe agent behavior. Agents with differing visual resolutions have access to differing numbers of possible visual observations. Lower visual resolution allows a lower total number of views than higher. A basis was chosen with visual resolution 8 with 132 unique views. The corresponding mean positions and orientations standardized the set of visual observations for every agent. The absolute value was taken for the resulting 132 actions, as agents learned to prefer clockwise or counterclockwise turning, and this break in symmetry is irrelevant to the classification of navigational strategy.
The results of two popular nonlinear dimensionality reduction algorithms, t-SNE and UMAP, both demonstrate separation aligning with our classification (S10 Fig). The direct pathing class distinctly separates from the two angle-based classes, and the indirect sequential and biased diffusive stretch along a gradient in another cluster. The clear distinction separating distance-based and angle-based navigational strategies, intuitively aligns with expectations. The graded cluster of the two angle-based classes suggests a continuum rather than separation, which may perhaps be more apparent when measures other than decorrelation time and directional entropy are taken into account. Despite such possibility, both decorrelation time and directional entropy appear to align with the separation and gradient of the two dimensionality reduction procedures. Though the former more smoothly aligns, especially considering the DP/BD hybrid agents.
Prior to t-SNE or UMAP, the data was passed through PCA to identify the 50 most principal components, a practice cited to reduce noise and high dimensionality. In practice, it did not appear to give a significantly different result.
Gaussian mixtures were applied to the t-SNE and UMAP results, demonstrating that unsupervised classification draws boundaries in a similar way to our metric-based procedure (S10 Fig).
@@ 5 Acknowledgements
P.G. acknowledges support by the Elsa-Neumann-Stipendium by the state of Berlin under Nachwuchsförderungsgesetz (NaFöG) - application number H75014. P.R. acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
All code and data used for this paper is publically available at https://github.com/pgovoni21/vis-nav-abm.
6 Supporting information
Fig S1: Evolutionary performance with differing parameters. Left: median performance of 40 training runs, dashed lines indicate median of validation tests. Right: validation test distribution (remaining distance to patch not included), lines indicate median, same color as legend in left plot. Runs used in top figure reflect data used in main text (perfect: theoretical lower bound). Unless otherwise noted, parameters are as follows: Distance scaling factor ($\sigma$): 0; Visual resolution ($\upsilon$): 8; Field of vision: 0.4; CNN output size: 4; Multilayer perceptron (MLP) size: 1x2.
(a) Indirect Sequential
(b) Biased Diffusive
(c) Direct Pathing
Fig S2: Extended figure from Fig 2 top row trajectories. Top/middle rows: $\upsilon=8$. Bottom row: $\upsilon=16$. See S3 Fig for corresponding correlations.
Fig S3: Extended figure from Fig 2 bottom row correlations. Top/middle rows: $\upsilon=8$. Bottom row: $\upsilon=16$. See S2 Fig for corresponding trajectories.
Fig S4: Perturbations to a trained agent (perception, environment, action). Top row: perturbation of field of vision. Middle row: perturbation of NW (top left) corner. Bottom row: perturbation of agent speed. Middle column: unperturbed agent. Red lines: visual guides.
Fig S5: Dual corner raycast as elliptical decision manifold. 1st row: ideal locations for double detection of the top-left/bottom-left corners, for differing field of vision parameters, where the agents in (b) illustrate example agent positions and orientations along one of the available elliptical arcs. As the agent is rotated, visual observation remains the same, allowing a consistent action to be tied across the manifold. 2nd row: trajectory maps with points plotted for simultaneous dual corner raycasts. 3rd row: absolute value of turning speed, normalized between 0 and 90 degrees per timestep. 4th row: same as 1st but for differing visual resolutions.
(a) Indirect Sequential
(b) Biased Diffusive
(c) Direct Pathing
Fig S6: Mean action vector fields. Mean vectors are computed and plotted for agent action outputs at 64 unique orientations for each grid location. Mean vector orientation are plotted with colors following the colorbar in Fig S5. Agents used mirror those in Fig 2 (top row) & S2 Fig (bottom three). Black points signify global minima in mean vector magnitude.
(a) Indirect Sequential
(b) Biased Diffusive
(c) Direct Pathing
Fig S7: Effect of noise to the visual angle. Visual angle, as the median of an agent FOV, is perturbed by Gaussian noise with the two standard deviation of 0.05 and 0.1 radians. Same set of initializations/networks as in Fig 2 (top row) & S2 Fig (bottom three).
Fig S8: A collection of trajectory plots representing behavioral hybridization possible. $\sigma$ damped to 0.5, midway between Fig 2 A/B & C, at visual resolution 8, arranged as a double continuum of classes: from indirect sequential to direct pathing to biased diffusion.
Fig S9: Extended data for Fig 4. Top left: density-based heatmap, maximum in yellow normalized at 5. Top right: relative fitnesses (blue: maximum, yellow: minimum). Bottom row: unsupervised clustering (left: K-Means, right: gaussian mixture), cluster centers plotted as larger circles.
Fig S10: Classification via dimensionality reduction: t-SNE (left) & UMAP (right). Sets of actions associated with visual observations from 132 positions and orientations in the environment for the trained agents of each evolutionary run were passed through the two dimensionality reduction procedures. Results are plotted according to: 1st row: classification labels from the metric-based procedure as illustrated in Fig 4, 2nd row: decorrelation time, 3rd row: directional entropy 4th row: Gaussian mixture models were fit to the results, redrawing similar class boundaries.